a man is 37 years old and his child is 8 years old. How many years ago was the product of their ages 96?
Answer by ankor@dixie-net.com(22740)

a man is 37 years old and his child is 8 years old.
How many years ago was the product of their ages 96?
:
Let y = no. of yrs ago that this was true
:
(37-y)(8-y) = 96
FOIL
296 - 37y - 8y + y^2 = 96
:
y^2 - 45y + 296 - 96 = 0
y^2 - 45y + 200 = 0
Factors to
(y-5)(y-40) = 0
Two solutions
y = 40
and
y = 5 yrs, is the only reasonable answer
:
5 yrs ago: 32 * 3 = 96